OpenStudy (anonymous):

The equation d = m/v can be used to calculate the density, d, of an object with mass, m, and volume, V. Which is an equivalent equation solved for V?

OpenStudy (anonymous):

I think it's m/d= V but then I think m-d = V is also right, please help!

OpenStudy (mathstudent55):

\(d = \dfrac{m}{v} \) First, multiply both sides by v \(vd = m\) Now divide both sides by d \(v = \dfrac{m}{d} \)

OpenStudy (mathstudent55):

You are correct.